Grade 12IIT JEE Entrance Exam

A block is resting on a horizontal plate in the xy plane and the coefficient of friction between block and plate is μ. The plate begins to move with velocityv = bt2 in x-direction. At what time will the block start sliding on the plate?

yash
11 Years agoGrade 12

5 Answers

PRATYUSH MISHRA
ApprovedApproved Tutor Answer11 Years ago
Hi yash!
as the plate begins to move frictional force acts on the block along the direction of velocity to prevent relative motion.
when the force exceed the maximum static friction the block will move.
a=dv/dt=2bt and the force by plate on block=ma(where m is the mass of block)
so when ma exceeds μmg then the block begins to slide
so ma>=μmg so t>=μg/2b
so at t=μg/2b the block slides
 
 
 
yash
11 Years ago
thanks 
 
yash
11 Years ago
are you studying in an iit 
PRATYUSH MISHRA
11 Years ago
no not yet yash I am a student of class 11
sajid
11 Years ago
as the plate begins to move frictional force acts on the block along the direction of velocity to prevent relative motion.
when the force exceed the maximum static friction the block will move.
a=dv/dt=2bt and the force by plate on block=ma(where m is the mass of block)
so when ma exceeds μmg then the block begins to slide
so ma>=μmg so t>=μg/2b
so at t=μg/2b the block slides